if the square root of a right triangle's hypotenuse is 10 and one of the legs is 28. What is the length of the last leg of the right triangle?

Answer by DrBeeee(684)

Using the usual notation of a right triangle the sides are a nd b and the hypotenuse is c, we have
(1) c^2 = a^2 + b^2
We are given that c= 100 and b = 28.
Using (1) we have
(2) a^2 = c^2 - b^2 or
(3) a^2 = 100^2 - 28^2 or
(4) a^2 = 10000 - 784 or
(5) a^2 = 9216 or
(6) a = sqrt(9216) or
(7) a = 96
Let's use (1) to check the answer.
Is (100^2 = 96^2 + 28^2)?
Is (10000 = 9216 + 784)?
Is (10000 = 10000)? Yes
Answer: The last leg of the right triangle is 96.